Castro25:

Titus works at a hotel. Part of his job is to keep the complimentary pitcher of water at least half full and always with ice. When he starts his shift, the water level shows 4 gallons, or 128 cups of water. As the shift progresses, he records the level of the water every 10 minutes. After 2 hours, he uses a regression calculator to compute an equation for the decrease in water. His equation is W –0.414t + 129.549, where t is the number of minutes and W is the level of water. According to the equation, after about how many minutes would the water level be less than or equal to 64 cups?

LilChris:

and what are the options choices

Castro25:

150 minutes 160 minutes 170 minutes 180 minutes

LilChris:

what do u think it is

AngeI:

W-.0414T+129.549 This is the equation correct?

LilChris:

@Ultrilliam

LilChris:

@umm

LilChris:

@Hero

Eiwoh2:

Angel is helping, Chris. No need to tag others.

AngeI:

-.0414(160)+129.549=64

AngeI:

Subtract 129.549 from both sides -.0414=64-129.549 -.4014=-65.549

AngeI:

Divide both sides by -.04014 getting 158.33 which is about 160

Castro25:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Castro25 thank you \(\color{#0cbb34}{\text{End of Quote}}\)

LilChris:

make sure when you post a question u always provide the answer choices, good job

Castro25:

thank you

AngeI:

No problem c: